Indicate how concentrated H2SO4 for which is 18 F may be diluted to produce 300 ml of 3 F H2SO4-


1
Expert's answer
2021-03-02T01:59:06-0500

As the initial solution is diluted, the volume-concentration is:

C2/C1 = V1/V2

where C1 - initial concentration, C2 - final concentration, V1 - volume of the initial solution, V2 - volume of the final solution.

From here:

V1 = V2 × C2 / C1 = 300 ml × 3 F / 18 F = 50 ml

Volume of water equals 300 ml - 50 ml = 250 ml

As a result, take 250 ml of water and add 50 ml of 18 F acid to obtain 300 ml of 3 F acid solution.